# Ensemble Kalman Filtering-Aided Variational Inference for Gaussian Process State-Space Models

### 23 Dec 2023 v2

@@ III Problem Statement
To overcome the intractability of the marginal distribution $p({\vec{\mathbf{y}}}|\bm{\theta})$, variational GPSSMs involve constructing an ELBO, denoted by $\mathcal{L}$, to the logarithm of the marginal likelihood, $\log p({\vec{\mathbf{y}}}|{{\bm{\theta}}})$. Specifically, the ELBO is constructed such that the difference between $\log p({\vec{\mathbf{y}}}|{{\bm{\theta}}})$ and $\mathcal{L}$ is equal to the Kullback-Leibler (KL) divergence between the variational approximation, $q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})$, and the true posterior, $p({\vec{\mathbf{x}}},{\vec{\mathbf{f}}}|{\vec{\mathbf{y}}})$, i.e.,
@@
Detailed derivations can be found in Supplement D-A. Maximizing $\mathcal{L}$ w.r.t. $\bm{\theta}$ fine-tunes the model parameters to improve the fit with the observed data, reflecting the model learning process; while maximizing $\mathcal{L}$ w.r.t. the variational distribution $q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})$ is equivalent to minimizing the KL divergence, $\operatorname{KL}[q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})\|p({\vec{\mathbf{x}}},{\vec{\mathbf{f}}}|{\vec{\mathbf{y}}})]$. That is, it enhances the quality of the variational distribution approximation, bringing it closer to the underlying posterior distribution, $p({\vec{\mathbf{x}}},{\vec{\mathbf{f}}}|{\vec{\mathbf{y}}})$, which corresponds to the model inference process [41]. The capacity of $q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})$ to approximate $p({\vec{\mathbf{x}}},{\vec{\mathbf{f}}}|{\vec{\mathbf{y}}})$ is thus of crucial importance [42].
@@
where $q({\vec{\mathbf{f}}})$ represents the variational distribution of the GP and $q({\mathbf{x}}_{0})\prod_{t=1}^{T}q({\mathbf{x}}_{t}|{\mathbf{f}}_{t})$ corresponds to the variational distribution of the latent states [43, 39]. It is noteworthy that the factorization of the variational distribution presented in Eq. (8) is recognized as a NMF approximation within the GPSSM literature (see formal definition in Appendix C) [30], because it explicitly establishes the dependencies between the latent states and the GP transition function values, as manifested by the terms $\prod_{t=1}^{T}q({\mathbf{x}}_{t}|{\mathbf{f}}_{t})$.
The methods closely related to this paper are probably [29] and [30]. In [29], variational distribution $q({\mathbf{x}}_{t}|{\mathbf{f}}_{t})$ is set equal to the prior distribution $p({\mathbf{x}}_{t}|{\mathbf{f}}_{t})$, i.e., $q({\mathbf{x}}_{t}|{\mathbf{f}}_{t})=p({\mathbf{x}}_{t}|{\mathbf{f}}_{t})$. But this choice results in no filtering or smoothing effect on latent states, although there is a dependence between latent states ${\vec{\mathbf{x}}}$ and transition function values ${\vec{\mathbf{f}}}$. Instead, the work in [30] assumes a parametric Markov-structured Gaussian variational distribution over the temporal states, i.e.,
@@
where $\{\bm{A}_{t},\bm{b}_{t},\bm{S}_{t}\}_{t=1}^{T}$ are free variational parameters. This choice, however, introduces a significant drawback: the number of variational parameters grows linearly with the length of the time series. Although this issue can be partially addressed by incorporating an inference network, such as a bidirectional recurrent neural network [28, 43], to learn the variational parameters and make the variational distribution learning with a constant number of parameters, fine-tuning these variational parameters in the inference network still requires substantial effort. Consequently, despite the flexibility of the variational distribution described in Eq. (9) allowing for approximation of the true posterior distribution, its empirical performance often falls short of its theoretical expressive capacity [32]. Moreover, it is worth noting that the inference network commonly takes the input of the observations, ${\vec{\mathbf{y}}}$, rendering the existing inference network-based methods predominantly trained in an offline manner, thus posing challenges when attempting to adapt them for online applications. To address these limitations, this paper proposes a novel approach that combines well-established EnKF with modern automatic differentiation tools [44] for variational inference in GPSSMs. The main objective is to alleviate the model and computational complexity and develop a more interpretable approach (e.g., stepping away from the potential black-box inference networks).
@@ IV-A Ensemble Kalman Filter (EnKF)
Given a SSM, see Eq. (II-B), the EnKF sequentially approximates the filtering distributions using $N\in\mathbb{N}$ equally weighted particles [45]. Specifically, at the prediction step, EnKF first samples particles, ${\mathbf{x}}_{t-1}^{1:N}\triangleq\{{\mathbf{x}}_{t-1}^{n}\}_{n=1}^{N}$, from the filtering distribution, $p({\mathbf{x}}_{t-1}|{\mathbf{y}}_{1:t-1})$, at time $t-1$. Then, the sampled particles are propagated using the state transition function $f(\cdot)$, i.e.,
@@
Thus, the filtering distribution, $p({\mathbf{x}}_{t}|{\mathbf{y}}_{1:t})\!=\!{\cal N}({\mathbf{x}}_{t}|\mathbf{m}_{t},\mathbf{P}_{t}),$ can be obtained using conditional Gaussian identity at the update step, where
@@
In the following subsections, we will describe our novel integration of the variational inference framework with the EnKF [45], which leads to a reduction in the number of variational parameters and ultimately enhances the learning and inference performance of GPSSMs.
@@ IV-B1 Sparse GPSSMs
Before introducing the utilization of EnKF, we introduce the sparse GP [34], a widely utilized technique in various GP variational approximations. This method serves as a scalable approach to model the corresponding GP component within the approximate posterior distribution (see Eq. (8)), thereby guaranteeing inherent scalability in GPSSM. The main idea of the sparse GPs is to introduce a small set of inducing points $\vec{{\mathbf{z}}}\triangleq\{{\mathbf{z}}_{i}\}_{i=1}^{M}$ and $\vec{{\mathbf{u}}}\triangleq\{{\mathbf{u}}_{i}\}_{i=1}^{M}$, $M\!\ll\!T$, to serve as the surrogate of the associated GPs, where the inducing inputs, ${\mathbf{z}}_{i}\!\in\!\mathbb{R}^{d_{x}},\forall i$, are placed in the same space as the latent states ${\mathbf{x}}_{t}$, while the corresponding inducing outputs ${\mathbf{u}}_{i}\!=\!f({\mathbf{z}}_{i})$ follow the same GP prior as ${\vec{\mathbf{f}}}$. With the augmentation of the inducing points, the joint distribution of the GPSSM becomes
@@ IV-B2 ELBO for sparse GPSSM
Before elucidating the methodology of employing EnKF to eliminate the parameterization for the variational distribution $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})$, let us first undertake the derivation of the ELBO induced by Eq. (16) and Eq. (19). After performing certain algebraic calculations, the ELBO is obtained as
Remark 1.
Each component of the ELBO in Eq. (21) is interpretable. More specifically,
@@ IV-B3 EnVI algorithm
We next proceed to outline the approximate evaluation of the ELBO in Eq. (21) using EnKF. The main idea revolves around utilizing the filtering distribution acquired from EnKF as a proxy for the corresponding variational distribution, acting as an approximation for the underlying posterior distribution, $p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\vec{\mathbf{y}}})$.
Building upon the EnKF outlined in Section IV-A and assuming that at time step $t-1$, we have acquired the variational distribution, $q({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}})\triangleq p({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})={\cal N}({\mathbf{x}}_{t-1}|\mathbf{m}_{t-1},\mathbf{P}_{t-1})$, we employ the GP transition, presented in Eq. (17), to perform the prediction step and generate $N$ predicted samples $\bar{{\mathbf{x}}}_{t}^{1:N}$, i.e.,
@@
where $\{{\mathbf{x}}_{t-1}^{n}\}_{n=1}^{N}$ are $N$ equally weighted particles generated from $q({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}})$. With the predicted samples $\bar{{\mathbf{x}}}_{t}^{1:N}$, we thus can approximate the prediction distribution as a Gaussian:
@@
and obtain the set of $N$ updated samples ${\mathbf{x}}_{t}^{1:N}$. Note that in the context of GPSSMs, both the prediction and update steps rely on conditioning on the inducing points, which act as a surrogate for the transition function.
@@
and both approximations $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}})$ and $q({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}})$ being Gaussian, determining the functional form of a common posterior transition $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})=p({\mathbf{x}}_{t}|{\mathbf{x}}_{t-1},{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t}),\forall t$, remains a challenging task. Therefore, alternative approximation methods are required to efficiently evaluate the ELBO. One alternative manner is to simply let $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})=p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})$, we then end up with the same ELBO as in [29], where no filtering or smoothing effect is feasible. To circumvent this limitation, we examine the difference between term 1 and term 4 in Eq. (21) and propose our EnKF-aided variational approximation. The main result is summarized in the following proposition.
@@
Under the approximations that, $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}})=p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t}),\forall t$, and, $p({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t})\approx p({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})$, the difference between term 1 and term 4 in Eq. (21) approximately corresponds to the expected log-likelihood, i.e.,
[…]
@@
Proposition 1 establishes that computing the expected log-likelihood yields an approximation to the difference between term 1 and term 4, thus the ELBO in Eq. (21) can be approximately evaluated as:
$\!-\operatorname{KL}[q({\vec{\mathbf{u}}})\|p({\vec{\mathbf{u}}})],$
where the log-likelihood, $\log p({\mathbf{y}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})$, at each step $t$ can be analytically evaluated by leveraging the Gaussian prediction distribution, see Eq. (23), and the linear emission model, i.e.,
and the KL divergence terms can be analytically computed in closed form [11], due to the Gaussian nature of the prior and variational distributions. Together with all, we can first sample ${\vec{\mathbf{u}}}\!\sim\!q({\vec{\mathbf{u}}})$ and numerically get an unbiased evaluation of $\mathcal{L}$. Thanks to the reparametrization trick [48], we then can apply gradient-based methods [49] to maximize the lower bound $\mathcal{L}$ w.r.t. the model parameters ${{\bm{\theta}}}=[\bm{\theta}_{gp},\mathbf{Q},\mathbf{R}]$ and the variational parameters $\bm{\zeta}=[\mathbf{m}_{0},\mathbf{L}_{0},\mathbf{m},\mathbf{S},{\vec{\mathbf{z}}}]$. The pseudo code for implementing the EnKF-aided variational learning and inference algorithm, termed as EnVI, is summarized in Algorithm 1.
@@
The analytical evaluation of the ELBO, $\mathcal{L}$, as illustrated in Eq. (27), circumvents the explicit evaluation of the first and fourth terms in Eq. (21), and sidesteps the computational challenges posed by the parameterized $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})$. Consequently, EnVI substantially can improve the efficiency of model learning and inference. In addition, maximizing the ELBO in Eq. (27) can be interpreted as follows: The objective is to optimize the model parameters and variational parameters such that the GPSSM can fit the observed data well at each step (indicated by the first term); simultaneously, the KL regularization terms impose constraints to prevent model overfitting (indicated by the second and third terms).

### 08 Jan 2024 v3

@@ III Problem Statement
To overcome the intractability of the marginal distribution $p({\vec{\mathbf{y}}}|\bm{\theta})$, variational GPSSMs involve constructing a model evidence lower bound (ELBO) to the logarithm of the marginal likelihood. Specifically, the ELBO, denoted by $\mathcal{L}$, is constructed such that the difference between $\log p({\vec{\mathbf{y}}}|{{\bm{\theta}}})$ and $\mathcal{L}$ is equal to the Kullback-Leibler (KL) divergence between the variational approximation, $q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})$, and the true posterior, $p({\vec{\mathbf{x}}},{\vec{\mathbf{f}}}|{\vec{\mathbf{y}}})$, i.e.,
@@
Detailed derivations can be found in Supplement D-A. Maximizing $\mathcal{L}$ w.r.t. $\bm{\theta}$ fine-tunes the model parameters to fit the observed data in the model learning process; while maximizing $\mathcal{L}$ w.r.t. the variational distribution $q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})$ is equivalent to minimizing the KL divergence, $\operatorname{KL}[q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})\|p({\vec{\mathbf{x}}},{\vec{\mathbf{f}}}|{\vec{\mathbf{y}}})]$. That is, it enhances the quality of the variational distribution approximation, bringing it closer to the underlying posterior distribution, $p({\vec{\mathbf{x}}},{\vec{\mathbf{f}}}|{\vec{\mathbf{y}}})$, which corresponds to the model inference process [37]. The capacity of $q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})$ to approximate $p({\vec{\mathbf{x}}},{\vec{\mathbf{f}}}|{\vec{\mathbf{y}}})$ is thus of crucial importance [38].
@@
where $q({\vec{\mathbf{f}}})$ represents the variational distribution of the GP and $q({\mathbf{x}}_{0})\prod_{t=1}^{T}q({\mathbf{x}}_{t}|{\mathbf{f}}_{t})$ corresponds to the variational distribution of the latent states [39, 35]. It is noteworthy that the factorization of the variational distribution presented in Eq. (8) is recognized as an NMF approximation within the GPSSM literature (see formal definition in Appendix C) [28], because it explicitly establishes the dependencies between the latent states and the GP transition function values, as manifested by the terms $\prod_{t=1}^{T}q({\mathbf{x}}_{t}|{\mathbf{f}}_{t})$.
The works most closely related to this paper are probably [27] and [28]. In [27], variational distribution $q({\mathbf{x}}_{t}|{\mathbf{f}}_{t})$ is subjectively set equal to the prior distribution $p({\mathbf{x}}_{t}|{\mathbf{f}}_{t})$, i.e., $q({\mathbf{x}}_{t}|{\mathbf{f}}_{t})=p({\mathbf{x}}_{t}|{\mathbf{f}}_{t})$. But using the prior distribution as an approximation results in no filtering or smoothing effect on latent states, although there is a dependence between the latent states ${\vec{\mathbf{x}}}$ and the transition function values ${\vec{\mathbf{f}}}$. Instead, the work in [28] assumes a parametric Markov-structured Gaussian variational distribution over the temporal states, i.e.,
@@
where $\{\bm{A}_{t},\bm{b}_{t},\bm{S}_{t}\}_{t=1}^{T}$ are free variational parameters. This choice, however, introduces a significant drawback: the number of variational parameters grows linearly with the length of the time series. Although this issue can be partially addressed by incorporating an inference network, such as a bidirectional recurrent neural network [26, 39], to learn the variational parameters and make the variational distribution learning with a constant model complexity, fine-tuning these variational parameters of the inference network still requires substantial effort. Consequently, despite the flexibility of the variational distribution described in Eq. (9) allowing for approximation of the true posterior distribution, its empirical performance often falls short of its theoretical expressive capacity [30]. Moreover, it is noteworthy that the inference network commonly takes the input of the observations, ${\vec{\mathbf{y}}}$, rendering the existing inference network-based methods predominantly trained in an offline manner, thus posing challenges when attempting to adapt them for online applications. To address these limitations, we step away from the heavy parameterization (e.g., the black-box inference networks) and propose a novel interpretable EnKF-aided algorithm for variational inference in GPSSMs. Our algorithm can exploit the dependencies between ${\vec{\mathbf{x}}}$ and ${\vec{\mathbf{f}}}$ and alleviate the model and computational complexities while being easily extended to an online learning algorithm.
@@ IV-A Ensemble Kalman Filter (EnKF)
The EnKF is a Monte Carlo-based method that excels in handling nonlinear systems compared to the KF and EKF [40]. Given a SSM, see Eq. (II-B), the EnKF sequentially approximates the filtering distributions using $N\!\in\!\mathbb{N}$ equally weighted particles [40]. Specifically, at the prediction step, EnKF first samples particles, ${\mathbf{x}}_{t-1}^{1:N}\triangleq\{{\mathbf{x}}_{t-1}^{n}\}_{n=1}^{N}$, from the filtering distribution, $p({\mathbf{x}}_{t-1}|{\mathbf{y}}_{1:t-1})$, at time $t-1$. Then, the sampled particles are propagated using the state transition function $f(\cdot)$, i.e.,
@@
Thus, the filtering distribution, $p({\mathbf{x}}_{t}|{\mathbf{y}}_{1:t})\!=\!{\cal N}({\mathbf{x}}_{t}|\mathbf{m}_{t},\mathbf{P}_{t}),$ can be obtained using the conditional Gaussian identity at the update step, where
@@
Remark 1.
Compared to another Monte Carlo-based method, PF, EnKF demonstrates computational efficiency, particularly in exceedingly high-dimensional spaces [41], though the PF may offer more flexibility in addressing highly non-Gaussian and nonlinear aspects. More importantly, the inherent differentiable nature of the EnKF (see Eqs. (10)–(14)) enables seamless integration with off-the-shelf automatic differentiation tools (e.g., PyTorch [42]) for model parameter optimization (see more discussions in Section IV-C). This stands in contrast to PF, where the discrete distribution resampling significantly poses challenges for the utilization of the reparameterization trick [43], resulting in significantly higher computational complexity for the parameters gradient computations [44, 45, 46, 47].
In the following subsections, we will describe our novel EnKF-aided NMF variational inference method, which significantly reduces the number of variational parameters and ultimately enhances the learning and inference performance of GPSSMs.
@@ IV-B1 Sparse GPSSMs
Before introducing the utilization of EnKF in the variational inference framework, let us first introduce the sparse GP [32], a widely utilized technique in various GP variational approximations. This method serves as a scalable approach to model the corresponding GP component within the approximate posterior distribution (see Eq. (8)), thereby guaranteeing inherent scalability in GPSSM. The main idea of the sparse GP is to introduce a small set of inducing points $\vec{{\mathbf{z}}}\triangleq\{{\mathbf{z}}_{i}\}_{i=1}^{M}$ and $\vec{{\mathbf{u}}}\triangleq\{{\mathbf{u}}_{i}\}_{i=1}^{M}$, $M\!\ll\!T$, to serve as the surrogate of the associated GP, where the inducing inputs, ${\mathbf{z}}_{i}\!\in\!\mathbb{R}^{d_{x}},\forall i$, are placed in the same space as the latent states ${\mathbf{x}}_{t}$, while the corresponding inducing outputs ${\mathbf{u}}_{i}\!=\!f({\mathbf{z}}_{i})$ follow the same GP prior as ${\vec{\mathbf{f}}}$. With the augmentation of the inducing points, the joint distribution of the GPSSM becomes
@@ IV-B2 ELBO for sparse GPSSM
Before elucidating the methodology of employing EnKF to eliminate the parameterization for the variational distribution $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})$, we first undertake the derivation of the ELBO, which is succinctly summarized as follows.
Theorem 1.
Upon the augmentation of induced points into the GPSSM (see Eq. (16)) and under the NMF assumption of the variational distribution (see Eq. (19)), the model evidence lower bound for joint learning and inference is:
$\!\!\mathcal{L}$ $=\underbrace{\mathbb{E}_{q({\vec{\mathbf{u}}},{\vec{\mathbf{x}}})}\left[\sum_{t=1}^{T}\log p({\mathbf{y}}_{t}|{\mathbf{x}}_{t})\right]}_{\mathrm{term~{}1}}\ -\ \underbrace{\operatorname{KL}[q({\mathbf{x}}_{0})\|p({\mathbf{x}}_{0})]}_{\mathrm{term~{}2}}$ (21)
$\quad\!-\!\underbrace{\operatorname{KL}[q({\vec{\mathbf{u}}})\|p({\vec{\mathbf{u}}})]}_{\mathrm{term~{}3}}-\underbrace{\mathbb{E}_{q({\vec{\mathbf{u}}},{\vec{\mathbf{x}}})}\left[\sum_{t=1}^{T}\log\frac{q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})}{p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})}\right]}_{\mathrm{term~{}4}}.$
Proof.
The proof can be found in Appendix A
An important observation lies in the interpretability of each component within the ELBO in Eq. (21). Specifically, when maximizing the ELBO for learning and inference:
@@ IV-B3 EnVI algorithm
The NMF assumption applied to the variational distribution (Eq. (19)) typically results in an intractable evaluation of the first and fourth terms within the ELBO [28]. We next proceed to outline the approximate evaluation of the ELBO using EnKF. The main idea revolves around utilizing the filtering distribution acquired from EnKF as a proxy for the corresponding variational distribution, acting as an approximation for the underlying posterior distribution, $p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\vec{\mathbf{y}}})$.
Building upon the EnKF outlined in Section IV-A and assuming that we have acquired the variational distribution, $q({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}})\triangleq p({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})\!=\!{\cal N}({\mathbf{x}}_{t-1}|\mathbf{m}_{t-1},\mathbf{P}_{t-1})$, at time step $t\!-\!1$, we can employ the GP transition, presented in Eq. (17), to perform the prediction step and generate $N$ predicted samples $\bar{{\mathbf{x}}}_{t}^{1:N}$, i.e.,
@@
where $\{{\mathbf{x}}_{t-1}^{n}\}_{n=1}^{N}$ are $N$ particles generated from $q({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}})$ with equal weights. With the predicted samples $\bar{{\mathbf{x}}}_{t}^{1:N}$, we thus can approximate the prediction distribution as a Gaussian:
@@
and obtain the set of $N$ updated samples ${\mathbf{x}}_{t}^{1:N}$. Note that in the context of GPSSMs, both the prediction and update steps are conditioned on the inducing points, which act as a surrogate for the transition function.
@@
and both approximations $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}})$ and $q({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}})$ being Gaussian, determining the functional form of a common posterior transition $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})=p({\mathbf{x}}_{t}|{\mathbf{x}}_{t-1},{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t}),\forall t$, remains a challenging task. Therefore, alternative approximation methods are required to efficiently evaluate the ELBO. One alternative manner is to simply let $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})=p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})$, we then end up with the same ELBO as in [27], where no filtering or smoothing effect is feasible. To circumvent this limitation, we examine the difference between term 1 and term 4 in Eq. (21) and propose our EnKF-aided variational lower bound approximation. The main result is summarized in the following proposition.
@@
Under the approximations of EnKF filtering (Eq. (24)) and $p({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t})\approx p({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})$, the ELBO presented in Theorem 1 can be reformulated as a summation over several simple terms:
$\!-\operatorname{KL}[q({\vec{\mathbf{u}}})\|p({\vec{\mathbf{u}}})],$
where the log-likelihood, $\log p({\mathbf{y}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})$ in the first term can be analytically evaluated, i.e.,
and the two KL divergence terms can also be analytically computed in closed form, due to the Gaussian nature of the prior and variational distributions [10].
@@
Based on the result in Proposition 1, we first utilize the reparametrization trick [43] to sample ${\vec{\mathbf{u}}}\!\sim\!q({\vec{\mathbf{u}}})$ and numerically get an unbiased evaluation of the expected log-likelihood, $\mathbb{E}_{q({\vec{\mathbf{u}}})}\!\left[\sum_{t=1}^{T}\log p({\mathbf{y}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})\right]$; then we apply gradient-based methods [50] to maximize the $\mathcal{L}$ in Eq. (26) w.r.t. the model parameters ${{\bm{\theta}}}=[\bm{\theta}_{gp},\mathbf{Q},\mathbf{R}]$ and the variational parameters $\bm{\zeta}=[\mathbf{m}_{0},\mathbf{L}_{0},\mathbf{m},\mathbf{S},{\vec{\mathbf{z}}}]$. Detailed routine for implementing the EnKF-aided variational learning and inference algorithm, termed as EnVI, is summarized in Algorithm 1.
@@
It is noteworthy that the newly derived ELBO, $\mathcal{L}$ circumvents the explicit evaluation of the first and fourth terms in Eq. (21), and sidesteps the computational challenges posed by the heavy parameterization of $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})$. Consequently, EnVI can substantially improve the efficiency of model learning and inference. In addition, maximizing the ELBO in Eq. (26) can be interpreted as follows: The objective is to optimize the model parameters and variational parameters such that the GPSSM can fit the observed data well at each step (indicated by the first term); simultaneously, the KL regularization terms impose constraints to prevent model overfitting (indicated by the second and third terms).